OpenStudy (anonymous):

The sequence {an} is defined by a0 = 1 and an+1 = 2an + 2 for n = 0, 1, 2, ... . What is the value of a3 ?

OpenStudy (anonymous):

a0=1 a1=2*1+2=4 a2=2*4+2=10 a3=2*10+2=22

OpenStudy (anonymous):

first if \(a_0=1\) then \(a_1=a_{0+1}=2a_0+2=2\times 1+2=4\)

OpenStudy (anonymous):

similarly we can find a4,a5,..

OpenStudy (anonymous):

my brain did not process that at all...hahah